# MDSGen: Fast and Efficient Masked Diffusion Temporal-Aware Transformers for Open-Domain Sound Generation

### v1

@@ 4.3 Main Results
Diff-Foley (Luo et al., 2023) * 10.55 56.67 6.49 93.92 0.36 860M 172×172$\times$172 ×
See and Hear (Xing et al., 2024) 21.35 19.23 6.94 58.14 18.25 1099M 220×220$\times$220 ×
FoleyCrafter (Zhang et al., 2024) 12.07 42.06 5.67 83.54 2.96 1252M 250×250$\times$250 ×
MDSGen-T (Ours) 500k 14.18 37.51 6.25 97.91 0.01 5M 1.0×1.0$\times$1.0 ×
MDSGen-S (Ours) 500k 12.92 44.38 6.29 98.32 0.02 33M $\times6$
MDSGen-B (Ours) 500k 11.19 52.77 6.27 98.55 0.05 131M 26.2×26.2$\times$26.2 ×
MDSGen-B (Ours) 800k 12.29 57.12 6.43 91.62 0.05 131M 26.2×26.2$\times$26.2 ×
@@
B. Flickr_SoundNet dataset. We use the models trained on VGGSound to test on SoundNet dataset to evaluate its generalization. First, through quantitative metrics in the sound source localization task with Flickr-SoundNet (Aytar et al., 2016) test set. Second, qualitatively compare the generated audio across different methods. As shown in Tab. 2, our method outperforms other methods on the CIoU metric (82.01%) closer to the ground truth (83.94%), while the AUC remains comparable (around 55.5%). It shows that our generated audio provided better-aligned features with the visual information to localize the sound source. Diff-Foley performs worst, indicating that it is more overfitting. We provide their visualizations in the Appendix.
@@ 5.1 Alignment Accuracy: A Confidence Score Perspective
We assess the enhancement of alignment accuracy in our method by analyzing confidence scores from the VGGSound test set, using the output of the sigmoid function from the trained classifier that predicts audio-video alignment. As illustrated in Fig. 4, FoleyCrafter (acc=83.54%) produces many low-confidence samples, indicating misalignment.
In contrast, Diff-Foley (93.9%) achieves a higher proportion of high-confidence scores. Remarkably, our method reaches an accuracy of 98.6%, significantly increasing the number of high-confidence samples and demonstrating superior audio-video alignment compared to other approaches.
@@ 5.2 Feature Dimensionality Reduction and Redundant Features in CAVP
Unlike Diff-Foley, which uses a U-Net-based Stable Diffusion model and incorporates all 32 video frame features for cross-attention, we found that reducing video features from 32 channels to a single channel significantly improves audio generation performance across all metrics (Tab. 3). Diff-Foley’s CAVP encodes video features at $32\times 512$, aligning with the audio’s 32-channel representation (also $32\times 512$) for latent-level alignment via contrastive learning. However, in the second stage, a mismatch arises as the VAE reduces audio dimensions to $4\times 16\times 64$ while video expands to $32\times 768$, leading to redundancy and inefficiencies. Reducing video dimensionality to $1\times 768$ representation acts as a bottleneck, simplifying learning and enhancing alignment, as supported by our results.
@@
5.3 Channel Selection from RGB for Mel-Spectrogram
@@
$R$ ($i$ = 0) 11.40 51.54 6.29 98.51
$G$ ($i$ = 1) 11.19 52.77 6.27 98.55
$B$ ($i$ = 2) 11.23 52.32 6.27 98.56
$\frac{1}{3}\sum_{r\in\{R,G,B\}}r$ 11.29 52.27 6.28 98.54
Notably, while the resulting mel-spectrograms (right figures) seem visually indistinguishable, the histograms highlight their differences. This emphasizes the importance of considering each channel’s statistics in generating the final mel-spectrogram $\widehat{\mathbf{X}}$, with further comparisons available in the Appendix.
@@ 5.4 Masking Diffusion Strategies
We explore various audio-masking methods in diffusion transformers (Fig. 3) and compare them to traditional image-based techniques like SAM used with ImageNet (Gao et al., 2023). Our findings reveal that audio data behaves differently from images (SAM), with incorporating temporal awareness (TAM) into the masking task significantly boosting performance across all metrics (Tab. 6), including a 4-point IS score increase from 48.66 to 52.77. Using a diffusion transformer without masking (DiT) leads to suboptimal results across all metrics, highlighting the critical role of masking in model learning.
@@ 5.5 Learned Weights of Reducer
We analyzed how our models allocate attention across video frames by visualizing their learned magnitude weights (Fig. 6). The upper figure shows that our model applies varying attention levels, with larger models exhibiting higher weights and more distinct differences. After softmax normalization (bottom figure), consistent trends are observed for various channels, though not all, with model B focusing more on channels 1, 2, and 32. These findings demonstrate that the Reducer effectively captures key features, selectively updating weights to prioritize relevant ones for audio generation.
@@ 5.7 Sampling Tools
Sampling Method. We employ DPM-Solver (Lu et al., 2022) with 25 steps for sampling during inference. We find that increasing from 25 to 50 steps with dynamic classifier-free guidance (Gao et al., 2023) can slightly improve the performance. We used CFG = 5 and power scaling $\alpha=0.01$ for the optimal setting. Classifier-Guidance (CG). We found that combining CFG and CG slightly improves alignment accuracy and KL, consistent with (Luo et al., 2023), but has no impact on other metrics (see Tab. 8), which differs from their findings. A thorough investigation of network architecture and additional datasets is needed to assess the complementary effects of CFG and CG fully, but this is beyond the scope of our work.
@@ 6 Conclusions
7 Limitations and Future Works
Our method offers fast inference, efficient parameter, and memory usage, and achieves top performance in alignment accuracy and IS score. However, it does have some limitations. First, like other diffusion models, it requires multiple sampling steps during generation. Second, while the VGGSound dataset is suitable for this study, it is not a large-scale dataset, which may not fully exploit the potential of our approach. Third, the current design handles a fixed video length of 8.2 seconds. We hope to incorporate recent advancements in single-step diffusion techniques into our framework in the future. Additionally, while video collection from online sources is increasingly feasible, it can still be time-consuming and storage-intensive, posing challenges for individual researchers. We leave further exploration in this area, along with addressing the fixed-length constraint, for future work.
@@ Appendix A Appendix
A.1 Training Cost
@@ A.2 More Setups
A.3 Comparing more Visualizations
@@
A.4 RGB channel selection for Mel-Spectrogram
@@
A.5 Mask Ratio Ablation
@@ A.6 More Directions
Our approach can inspire applications of masked diffusion models across various domains such as object detection (Vu et al., 2019), video question answering (Kim et al., 2020), image super-resolution (Niu et al., 2024a; b), and speech processing. These models can also be combined with self-supervised learning techniques (He et al., 2022; Pham et al., 2021; 2022b; 2023; 2022a) to enhance representation learning for diverse generative tasks. With the potential of diffusion transformers for conditional learning, we anticipate further exploration of their capabilities in speech processing (Jung et al., 2022; 2020; Trung & Yoo, 2019), data augmentation (Lee et al., 2020), VQA (Kim et al., 2020), visual detection (Vu et al., 2019), and super-resolution (Niu et al., 2023; 2024a; 2024b).

### v2

@@ 4.3 Main Results
Diff-Foley (Luo et al., 2023) * 4.71 10.55 56.67 6.49 93.92 0.36 860M 172×172$\times$172 ×
See and Hear (Xing et al., 2024) 5.55 21.35 19.23 6.94 58.14 18.25 1099M 220×220$\times$220 ×
FoleyCrafter (Zhang et al., 2024) 2.45 12.07 42.06 5.67 83.54 2.96 1252M 250×250$\times$250 ×
MDSGen-T (Ours) 500k 2.21 (3.69) 14.18 37.51 6.25 97.91 0.01 5M 1.0×1.0$\times$1.0 ×
MDSGen-S (Ours) 500k 1.66 (2.75) 12.92 44.38 6.29 98.32 0.02 33M $\times6$
MDSGen-B (Ours) 500k 1.34 (2.16) 11.19 52.77 6.27 98.55 0.05 131M 26.2×26.2$\times$26.2 ×
MDSGen-B (Ours) 800k - 12.29 57.12 6.43 91.62 0.05 131M 26.2×26.2$\times$26.2 ×
@@
B. Flickr_SoundNet dataset. We use the models trained on VGGSound to test on SoundNet dataset to evaluate its generalization. First, through quantitative metrics in the sound source localization task with Flickr-SoundNet (Aytar et al., 2016) test set. Second, qualitatively compare the generated audio across different methods. As shown in Tab. 2, our method outperforms other methods on the CIoU metric (82.01%) closer to the ground truth (83.94%), while the AUC remains comparable (around 55.5%). It shows that our generated audio provided better-aligned features with the visual information to localize the sound source. Diff-Foley performs worst, indicating that it is more overfitting. We provide their visualizations in the Appendix A.
@@ 5.1 Alignment Accuracy: A Confidence Score Perspective
We assess the enhancement of alignment accuracy in our method by analyzing confidence scores from the VGGSound test set, using the output of the sigmoid function from the trained classifier that predicts audio-video alignment. As illustrated in Fig. 4, FoleyCrafter (acc=83.54%) produces many low-confidence samples, indicating misalignment. In contrast, Diff-Foley (93.9%) achieves a higher proportion of high-confidence scores. Remarkably, our method reaches an accuracy of 98.6%, significantly increasing the number of high-confidence samples and demonstrating superior audio-video alignment compared to other approaches.
@@ 5.2 Feature Dimensionality Reduction and Redundant Features in CAVP
Unlike Diff-Foley, which uses a U-Net-based Stable Diffusion model and incorporates all 32 video frame features for cross-attention, we found that reducing video features from 32 channels to a single channel significantly improves audio generation performance across all metrics (Tab. 3). Diff-Foley’s CAVP encodes video features at $32\times 512$, aligning with the audio’s 32-channel representation (also $32\times 512$) for latent-level alignment via contrastive learning. However, in the second stage, a mismatch arises as the VAE reduces audio dimensions to $4\times 16\times 64$ while video expands to $32\times 768$, leading to redundancy and inefficiencies. Reducing video dimensionality to $1\times 768$ representation acts as a bottleneck, simplifying learning and enhancing alignment, as supported by our results in the Appendix A.2.
@@
5.3 Does Neural Vocoder Helps?
We conducted experiments by training a separate neural vocoder to convert mel-spectrograms back to waveforms. Using the publicly available HiFi-GAN code from GitHub, we trained the vocoder from scratch on the VGGSound dataset, achieving good convergence within three days. As shown in Tab. 5, our method achieves an FAD score of 4.3788 with the simple Griffin-Lim algorithm, outperforming both See-and-Hear (5.5547) and Diff-Foley (6.0810). When enhanced with the HiFi-GAN neural vocoder, our method further improves the FAD score to 2.1610, achieving state-of-the-art performance on this metric. This result not only highlights the superiority of our approach but also demonstrates its robustness when evaluated beyond the metrics reported in Tab. 1.
Table 5: FAD on VGGSound test set. Our method MDSGen-B using the simple Griffin-Lim outperforms the two methods and is state-of-the-art when equipped with a vocoder HifiGAN.
Method See-and-hear FoleyCrafter Diff-Foley Diff-Foley MDSGen-B MDSGen-B
(neural vocoder) (neural vocoder) (Griffin-Lim) (neural vocoder) (Griffin-Lim) (neural vocoder)
@@ 5.3 Does Neural Vocoder Helps?
5.4 Subjective Human Evaluation Tests
We conducted a human evaluation by generating 50 audio samples based on 50 videos for each method. Five participants were asked to evaluate each method. Participants were instructed to watch the videos and listen to the corresponding audio, rating each on a scale from 1 to 5 based on the following criteria: 1) Audio Quality (AQ): How good is the sound quality? and 2) Audio-Video Content Alignment (AV): How well does the sound match the video content? The mean opinion scores (MOS) for each metric (ranging from 1 to 5) are presented in Tab. 6. The results from our human evaluation demonstrate that the waveforms generated by our model outperform those of competing methods, receiving higher scores across the evaluation criteria. Participants consistently rated the audio quality and alignment with the video content more favorably for our generated waveforms, indicating superior perceptual performance.
Table 6: Human Evaluation (MOS). Our method MDSGen-B equipped with a vocoder HifiGAN. AQ: Audio Quality, AV: Audio-visual content relevance.
Method See-and-hear FoleyCrafter Diff-Foley MDSGen-B (Ours) Ground Truth
[…]
@@ 5.5 Masking Diffusion Strategies
We explore various audio-masking methods in diffusion transformers (Fig. 3) and compare them to traditional image-based techniques like SAM used with ImageNet (Gao et al., 2023). Our findings reveal that audio data behaves differently from images (SAM), with incorporating temporal awareness (TAM) into the masking task significantly boosting performance across all metrics (Tab. 7), including a 4-point IS score increase from 48.66 to 52.77. Using DiT without masking leads to suboptimal results across all metrics, highlighting the critical role of masking in model learning.
@@
TAM outperforms FAM, as masking in the temporal dimension typically yields better performance in audio generation. This is likely due to the stronger impact of temporal structure on coherence, perceptual quality, and the dependencies within audio sequences.
@@ 5.6 Learned Weights of Reducer
We analyzed how our models allocate attention across video frames by visualizing their learned magnitude weights (Fig. 5). The upper figure shows that our model applies varying attention levels, with larger models exhibiting higher weights and more distinct differences. After softmax normalization (bottom figure), consistent trends are observed for various channels, though not all, with model B focusing more on channels 1, 2, and 32. These findings demonstrate that the Reducer effectively captures key features, selectively updating weights to prioritize relevant ones for audio generation. More experiments of Reducer choices are available in the Appendix A.3.
@@ 5.8 Sampling Tools
Sampling Method. We employ DPM-Solver (Lu et al., 2022) with 25 steps for sampling during inference. We find that increasing from 25 to 50 steps with dynamic classifier-free guidance (Gao et al., 2023) can slightly improve the performance. We used CFG = 5 and power scaling $\alpha=0.01$ for the optimal setting. Classifier-Guidance (CG). We found that combining CFG and CG slightly improves alignment accuracy and KL, consistent with (Luo et al., 2023), but has no impact on other metrics (Tab. 9), which differs from their findings. A thorough investigation of network architecture and additional datasets is needed to assess the complementary effects of CFG and CG, which is beyond the scope of our work.
@@ 6 Conclusions
Limitations and Future Works
Our method offers fast inference, efficient parameter usage, and low memory consumption, while achieving top performance in alignment accuracy and IS score. However, there are some limitations. First, like other diffusion models, it requires multiple sampling steps during generation. Second, while the VGGSound dataset is suitable for this study, its size may not fully leverage the potential of our approach. Third, the current design is constrained to a fixed video length of 8.2 seconds. In the future, we aim to incorporate recent advancements in single-step diffusion techniques to address this limitation. Additionally, although video collection from online sources is becoming more feasible, it remains time-consuming and storage-intensive, which may be challenging for individual researchers. We plan to explore the potential of 1D VAEs for further improvement and to address the fixed-length constraint in future work.
Acknowledgment
This work was supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2021-II211381, Development of Causal AI through Video Understanding and Reinforcement Learning, and Its Applications to Real Environments) and (No. RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics).
@@ Appendix A Appendix
A.1 Reducer Details
We show the simple design of our Reducer that can help to obtain global information while retaining local information:
  • Input: Video feature $V\in\mathbb{R}^{32\times 512}$
  • Output: feature vector $v\in\mathbb{R}^{1\times 768}$
Fig. 7 illustrates the details of the proposed Reducer architecture with two layers: the fully connected layer captures the local information of the video, and the second layer (1x1 conv) extracts the global information. Specifically, after the initial layer with fully connected weights, each dimension component (position) from 1 to 768 in the output vector $u_{1}$ contains the whole vector $v_{1}$ (local information of video). Consequently, in the next layer, each component of the final vector $v\in\mathbb{R}^{1\times 768}$ captures both local and global information from the original 32x512 video information. This ensures that the final vector provides comprehensive information for the subsequent audio generation process. This lightweight vector significantly reduces the burden of the DiT process.
Figure 7: Reducer Architecture. It includes a linear layer Linear(512,768) and a 1x1 conv layer Conv1d(32,1) that helps to retain local information while reducing dimension.
A.2 Overfitting Phenomenon with Redundant Features
We observe that the model becomes quickly overfitted if using redundant features in Fig. 8. By contrast, our Reducer helps to mitigate such redundant features and we can see that the test accuracy remains quite a close gap with train accuracy from 100k steps to 500k steps.
Figure 8: Overfitting with Redundant Features. We see that the redundant features show a bigger gap of overfitting where the test and train accuracy gap becomes larger.
A.3 Reducer Choices
We conducted an ablation study on the choice of the Reducer using three approaches: 1) Naive average pooling, 2) Attention pooling, and 3) Learnable weights. For this study, we used the MDSGen-B model trained for 300k steps. The results, shown in Tab. 11, indicate that all pooling methods achieve competitive performance with comparable alignment accuracy (Align. Acc). However, Learnable Weights yield the highest inception score (IS) at 7 points and slightly outperform in terms of FID. Meanwhile, Attention Pooling achieves the best KL metric.
Table 11: Reducer Vector Pooling Choice. Model MDSGen-B is trained for 300k steps
[…]
Naive Average Pooling 12.5594 39.41122 6.1750 0.9848
Attention Pooling 12.2704 39.0696 6.0819 0.9847
Learnable Weight (default) 12.1534 46.3301 6.3066 0.9858
We hypothesize that Attention Pooling performs better on the KL metric because its adaptive weights focus on the most relevant features in the input, enabling a more precise reconstruction of the latent distribution and, consequently, better KL divergence minimization. On the other hand, the Learnable Weights method performs best on the inception score because it directly optimizes the contribution of each dimension, tailoring the representation for the final task. This flexibility allows the model to capture both global and local information more effectively, leading to improved perceptual quality as reflected in the IS metric.
Learnable weights can indeed be considered a form of adaptive weighting since the weights are optimized during training and dynamically adjusted based on the data and task requirements. The distinction lies in the mechanism:
  • Attention pooling calculates adaptive weights based on the input features themselves (using attention scores). This is data-dependent and can adapt to specific patterns in the input at each forward pass.
  • Learnable weights, on the other hand, are parameterized and optimized during training, making them adaptable over time but not directly dependent on the input features in real time.
So while both methods involve adaptivity, attention pooling adapts dynamically per input, whereas learnable weights are statically optimized across the dataset. Naive average pooling is less effective compared to attention pooling and learnable weights because it assigns equal importance to all input features, regardless of their relevance to the task. This uniform weighting lacks the ability to focus on critical features or filter out irrelevant ones, which can dilute the quality of the pooled representation.
A.4 Choice of Decoder Layers
We compared the number of decoder layers and showed that $N_{2}=4$ gives better results on multiple metrics compared to $N_{2}=2$ as in the below table:
Table 12: Choice of decoder layer. Model MDSGen-B is trained for 300k steps
[…]
$N_{2}=2$ 12.4602 47.9981 6.4344 0.9823
$N_{2}=4$ (default) 12.1534 46.3301 6.3066 0.9858
A.5 Training Cost
@@ A.6 More Setups
A.7 Comparing more Visualizations
@@
A.8 Channel Selection from RGB for Mel-Spectrogram
@@
[…]
$R$ ($i$ = 0) 11.40 51.54 6.29 98.51
$G$ ($i$ = 1) 11.19 52.77 6.27 98.55
$B$ ($i$ = 2) 11.23 52.32 6.27 98.56
$\frac{1}{3}\sum_{r\in\{R,G,B\}}r$ 11.29 52.27 6.28 98.54
Notably, while the resulting mel-spectrograms (right figures) seem visually indistinguishable, the histograms highlight their differences. This emphasizes the importance of considering each channel’s statistics in generating the final mel-spectrogram $\widehat{\mathbf{X}}$, with more comparisons in the Appendix.
A.9 Mask Ratio Ablation
@@ A.10 More Directions
Our approach can inspire applications of masked diffusion models across various domains such as object detection (Vu et al., 2019), video question answering (Kim et al., 2020), image super-resolution (Niu et al., 2024a; b), and speech processing. These models can also be combined with self-supervised learning techniques (He et al., 2022; Pham et al., 2021; 2022b; 2023; 2022a) to enhance representation learning for diverse generative tasks. With the potential of diffusion transformers for conditional learning, we anticipate further exploration of their capabilities in speech processing (Jung et al., 2022; 2020; Trung & Yoo, 2019), data augmentation (Lee et al., 2020), VQA (Kim et al., 2020), visual detection (Vu et al., 2019), and super-resolution (Niu et al., 2023; 2024a; 2024b). Our work demonstrates the potential of Masked Diffusion Transformers for audio generation while also highlighting their applicability to various other tasks, including image editing (Koo et al., 2024; Song et al., 2024; Yoon et al., 2024) and personalized image generation using flow-based methods (Pham et al., 2024; Lipman et al., 2022).
Appendix B VAE Choices for Melspectrogram Reconstruction
We conducted an experiment to assess the impact of the VAE in our overall framework. Specifically, we compare the often used Stable Diffusion VAE which is used in Diff-Foley (Luo et al., 2023) and Action2Sound (Chen et al., 2024) and compared with the case when we employ the AudioLDM VAE to replace Stable Diffusion VAE. The results are shown in the below tables. Several key findings are observed. Below, we provide a side-by-side comparison highlighting the performance differences between the image VAE and the AudioLDM VAE across key metrics such as FAD, IS, KL, and Alignment Accuracy, and efficiency.
Part 1. Performance Comparison. We find that although the AudioLDM VAE achieves better results in FAD and KL, with a comparable FID score, its performance in Alignment Accuracy and IS is not as strong as that of the image VAE. Our smallest model, MDSGen-T (5M), achieves a state-of-the-art FAD score of 2.21, surpassing existing video-to-audio methods (the FAD of FoleyCrafter is 2.45). The larger model, MDSGen-B (131M), further improves the FAD to 1.34. Note that, we reported FAD 2.16 of our MDSGen in the previous response from checkpoint 500k training steps with image VAE. The detailed results are summarized in Tab. 1 below (300k steps):
Table 16: Performance Comparison AudioVAE: Image VAE vs. AudioLDM VAE
[…]
MDSGen-B (131M) ImageVAE (Rombach et al., 2022) 12.15 46.41 6.31 98.58 2.36
AudioLDM VAE (Liu et al., 2023) 11.83 30.37 6.03 96.52 1.34
MDSGen-S (33M) ImageVAE (Rombach et al., 2022) 12.88 40.39 6.21 98.25 2.75
AudioLDM VAE (Liu et al., 2023) 13.14 26.05 5.83 95.63 1.66
MDSGen-Tiny (5M) ImageVAE (Rombach et al., 2022) 14.78 33.23 6.23 97.62 3.69
AudioLDM VAE (Liu et al., 2023) 15.48 21.99 5.58 95.53 2.21
Part 2. Efficiency Comparison. We observed a slight difference in efficiency between the two VAEs, primarily due to variations in input dimensions and output characteristics. For instance, the AudioLDM VAE produces 816 latent tokens, which is 3.1x longer than the 256 tokens generated by the image VAE. As a result, models using the AudioLDM VAE tend to be slower and require more memory. However, it’s important to note that all variants of our model remain considerably more efficient than existing baseline methods. The details are shown in Tab. 2 below:
Table 17: Efficiency Comparison: Image VAE vs. AudioLDM VAE
VAE MelSpec. Res. Latent #Tokens Models #Prams. (M)$\downarrow$ Infer. Time (s)$\downarrow$ #Memory (M)$\downarrow$
Image VAE $128\times 512$ $16\times 64$ 256 MDSGen-T 5M 0.01 1406
MDSGen-S 33M 0.02 1508
MDSGen-B 131M 0.05 2132
AudioLDM VAE $64\times 816$ $16\times 204$ 816 MDSGen-T 5M 0.058 3087
MDSGen-S 33M 0.114 5383
MDSGen-B 131M 0.243 10165
VAE Analysis
Insight 1. Compared to existing video-to-audio methods, both VAEs show that MDSGen achieves state-of-the-art performance in FAD (1.34–2.16) and Alignment Accuracy (96.5%–98.5%). While the Image VAE outperforms the AudioLDM VAE by about 2% in Alignment Accuracy, it falls behind in FAD (2.16 vs. 1.34), indicating that AudioLDM produces higher-quality audio. This discrepancy in alignment accuracy and IS may be attributed to several factors:
Alignment Classifier Training: The alignment classifier is trained on 128×512 spectrograms, whereas AudioLDM uses 64×816 mel-spectrograms, inherently favoring features from the Image VAE, which aligns more closely with its input format. Domain-Specific Design: AudioLDM is optimized for audio reconstruction, resulting in features less visually aligned with the mel-spectrograms expected by the classifier. Resolutions: Image VAEs operate at higher spatial resolutions (128×512) and preserve finer, visually coherent patterns, leading to higher IS scores, while AudioLDM’s lower resolution (64×816) focuses on acoustic fidelity rather than visual consistency. Overall, we hypothesize that the Image VAE performs better in IS and alignment accuracy due to its compatibility with visual-based metrics and the image-like nature of spectrograms. In contrast, Audio VAEs prioritize audio quality, creating a trade-off between perceptual fidelity and visual alignment in video-to-audio tasks.
Insight 2. Both configurations show that MDSGen is highly efficient, utilizing significantly fewer parameters than existing approaches while maintaining strong performance across key metrics, including FAD, FID, and audio-video alignment accuracy.
Insight 3. The AudioLDM VAE improves audio quality over the image VAE, but at the cost of slower inference and slightly lower audio-video content relevance. Overall, our results reveal a trade-off: integrating MDSGen with the image VAE achieves the highest Alignment Accuracy and faster inference, while the AudioLDM VAE delivers superior audio quality (FAD) with slower inference.
In summary, our masked diffusion model, MDSGen, is the first to support both image- and audio-based VAEs for the video-to-audio task, offering a well-balanced trade-off between alignment accuracy and high-quality audio, while maintaining exceptional efficiency in parameters, memory usage, and inference speed. In the revised PDF, we have highlighted this in Section 3.5: VAE Choice for Mel-Spectrogram, where we discuss replacing the channel selection used in the RGB VAE. Additionally, we have moved the RGB-related discussion to the Appendix and will include a detailed analysis demonstrating the compatibility of our model with both image and audio VAEs.